OpenStudy (anonymous):

The midpoint of UV is (5,-11). The coordinates of one endpoint are U(3,5). Find the coordinates of endpoint V. I need the answer in steps. Help me and I will fan you, and give you a medal!

OpenStudy (leozap1):

if we have the end point at (0,0) then the midpoint is easy to work with :) (3, 5) (5,–11) (x , y ) -3 -5 -3 -5 -3 -5 ----------------------- (0, 0) (2,-16) (x-3, y-5) since the midpoint is half way, double it to get to the end. (3, 5) (5,–11) (x , y ) -3 -5 -3 -5 -3 -5 ----------------------- (0, 0) (2,-16) (x-3, y-5) 4 -32 --------- (x-1 , y-37) we modified (x=3,y=5) to begin with so lets sub those values back in. (3-1, 5-37) = (2, -27)

OpenStudy (anonymous):

I am still a little confused on how to put this into steps.

OpenStudy (wolfe8):

I'd say first you sketch to know which point is on which side.  Now you know the point you want it to the right and down of the midpoint. Midpoint means that from it to either ends of the line is the same distance. You have one of the points, so you can calculate the distance in the x-axis and y-axis. Then you can simple add that distance to the midpoint to get the other point. Does that make sense?